### v1

@@ 5 Discussion and Conclusion
In this report, we introduce Uni-SMART for in-depth understanding of multimodal information within the scientific literature. Through rigorous quantitative evaluation, Uni-SMART demonstrates significant performance gain in interpreting and analyzing multimodal contents in scientific documents, such as tables, charts, molecular structures, and chemical reactions, compared with other leading competitors. The success of Uni-SMART lies in its innovative cyclic iterative process that continuously refines its multimodal understanding capabilities, leveraging a robust dataset and a combination of multimodal learning, supervised fine-tuning, user feedback, expert annotation, and data enhancement to achieve superior performance in scientific literature analysis.
@@
Appendix
Appendix A Experiment Setup
For the evaluation of tables, charts, and chemical reactions, we conducted tests by uploading articles containing such content in PDF format and then posing questions to the models. Given the varying abilities of different models to process PDFs, we employed distinct methods to provide the PDFs.
For GPT-4, we utilized the web-based ChatGPT4 interface. We directly uploaded the original PDF files to the chat interface and posed queries, thereby leveraging OpenAI’s built-in PDF processing capabilities.
For GPT-3.5, we processed the PDFs using PyPDF2 to convert them into plain text, which was then inputted into the models.
For Gemini, given its proficiency in concurrently processing text and images, we initially employed PyPDF2 to extract text from the PDFs. Subsequently, we utilized PyMuPDF to retrieve images within the documents. Both text and images were then relayed to the model.
For Uni-SMART, owing to its capability to directly read PDF files, we simply uploaded the original PDF files and posed questions to the model.

### v2

@@ 5 Discussion and Conclusion
In this report, we introduce Uni-SMART for in-depth understanding of multimodal information within the scientific literature. Through rigorous quantitative evaluation, Uni-SMART demonstrates significant performance gain in interpreting and analyzing multimodal contents in scientific documents, such as tables, charts, molecular structures, and chemical reactions, compared with other competitors. The success of Uni-SMART lies in its innovative cyclic iterative process that continuously refines its multimodal understanding capabilities, leveraging a robust dataset and a combination of multimodal learning, supervised fine-tuning, user feedback, expert annotation, and data enhancement to achieve superior performance in scientific literature analysis.